# Length-Aware DETR for Robust Moment Retrieval

### v1

@@ Abstract
Video Moment Retrieval (MR) aims to localize moments within a video based on a given natural language query. Given the prevalent use of platforms like YouTube for information retrieval, the demand for MR techniques is significantly growing. Recent DETR-based models have made notable advances in performance but still struggle with accurately localizing short moments. Through data analysis, we identified limited feature diversity in short moments, which motivated the development of MomentMix. MomentMix employs two augmentation strategies: ForegroundMix and BackgroundMix, each enhancing the feature representations of the foreground and background, respectively. Additionally, our analysis of prediction bias revealed that short moments particularly struggle with accurately predicting their center positions of moments. To address this, we propose a Length-Aware Decoder, which conditions length through a novel bipartite matching process. Our extensive studies demonstrate the efficacy of our length-aware approach, especially in localizing short moments, leading to improved overall performance. Our method surpasses state-of-the-art DETR-based methods on benchmark datasets, achieving the highest R1 and mAP on QVHighlights and the highest R1@0.7 on TACoS and Charades-STA (such as a 2.46% gain in R1@0.7 and a 2.57% gain in mAP average for QVHighlights). The code is available at https://github.com/sjpark5800/LA-DETR.
@@ 1 Introduction
As vast amounts of video content are created and shared on the internet daily [7], the need for effective filtering has become more critical. Text-based search algorithms [3] have emerged as one of the most effective solutions, enabling rapid and accurate retrieval of videos that match user queries. To improve user experience and search efficiency, moment retrieval (MR) [1, 11] has gained significant attention. MR identifies the specific moments within a video that best align with a given query. Specifically, this task involves localizing the start and end points in the video relevant to the textual query, offering a more fine-grained understanding of video content.
@@
Table 1: Mean Average Precision (mAP) of previous moment retrieval (MR) DETR-based methods on QVHighlights val set by the lengths of the moment. Notice that the mAP drops significantly in capturing short-moment, where MR can be best utilized.
Method Short Middle Long All
${}_{\textrm{(seconds)}}$ (000,010] (010,030] (030,150] (000,150]
QD-DETR (CVPR 2023) 8.34 43.54 47.80 41.24
TR-DETR (AAAI 2024) 9.93 46.91 51.71 45.09
UVCOM (CVPR 2024) 12.65 49.04 49.39 45.80
For the MR task, previous approaches leveraging DETR for its efficiency and flexibility [19, 26, 32, 44, 38] have proposed methods to improve video-text feature representation, achieving impressive performance. However, our empirical findings indicate that these DETR-based models suffer from a significant drop in performance when handling short moments as highlighted in Table 1. For example, UVCOM [38] shows an average mAP of 49.04 for middle-length moments (10–30 seconds), while achieving only 12.65 for short moments (less than 10 seconds), revealing a substantial gap. As illustrated in Figure 1, moments are uniformly distributed across different lengths. However, due to the inherently short duration of short moments, they appear more frequently within a single video sample. Performance metrics are calculated by averaging the performance across all moments within a sample and then averaging these sample performances to obtain the overall performance. Therefore, the overall performance does not sufficiently reveal the performance on short moments, leading to a lack of attention to the challenges associated with short moment retrieval.
@@
Retrieving short moments within videos is a crucial task because videos often contain a significant amount of redundant or irrelevant information, while essential content is frequently condensed in short moments. This aligns with the importance of MR, where improving the accuracy of short-moment retrieval enables the precise extraction of the most relevant information. Such improvements can significantly reduce the time and effort required for video exploration. For example, highlights in sports and news, as well as key scenes in movies and dramas, often involves short moments. This emphasizes the importance of accurately retrieving short moments in practical scenarios.
Figure 1: In QVHighlights dataset, a single video can contain multiple moments of varying lengths. The yellow bars represent the count of each moment length across all videos. In contrast, the blue bars categorize a video sample by a specific length if it includes at least one moment of that length.
@@
In this study, we analyzed the challenges associated with existing methods in short-moment retrieval from both data and model perspectives. From a data perspective, we examined the feature distribution of short moments compared to other moments. As shown in Figure 2, the features of short moments tend to be more concentrated around the mean feature, with 42.9% of samples falling within one standard deviation. In contrast, for other moments, 26.6% of samples fall within one standard deviation of the mean. These observations show that short moments exhibit relatively simple and less diverse feature distributions. On the model side, we analyzed the trends in prediction accuracy by breaking down the model’s final output into center and length components, as illustrated in Figure 3. Although conceptually, moments are defined by (start, end), existing models predict them in the format of (center, length). Interestingly, we found that the accuracy of center predictions for short moments was significantly lower than for other moment types.
@@
Based on the above analysis, we propose a new DETR-based MR framework that addresses the performance degradation when retrieving short moments. Our framework consists of two novel techniques: a data augmentation technique called MomentMix and a length-aware decoder (LAD). Through data analysis, we identified a key limitation in the feature diversity of short moments. It prompts us to design MomentMix, which comprises two types of mix-based data augmentation strategies: ForegroundMix and BackgroundMix. In a video sample, we define the temporal moments relevant to the text query as the foreground and the unrelated moments as the background. ForegroundMix enhances the diversity of foreground features by combining rich foreground elements from different moments to create new short foregrounds. These newly created short foregrounds enable the model to detect short moments more robustly. Conversely, BackgroundMix utilizes portions of other videos as backgrounds, forming varied foreground-background combinations. This enables the model to better differentiate between foreground and background, ultimately improving the quality of video features across all moment lengths, not just short ones.
@@
Our analysis of model outputs revealed that center prediction errors significantly contribute to performance drops in short-moment retrieval. To address this, we introduce a length-aware decoder, which structures the decoder to focus more effectively on center prediction by conditioning it on length. Specifically, we predefine length classes (e.g., short, middle, long) and uniformly assign each decoder query to these length-specific classes. Additionally, we modify the bipartite matching process so that queries are matched with ground-truth moments within the same length class. This approach improves the accuracy of both center and length predictions for short moments.
In this work, we tackled the significant challenge of short-moment retrieval by proposing MomentMix augmentation and Length-Aware Decoder. Our approach significantly improved performance across various datasets, including QVHighlights, TACoS, and Charades-STA. Our contributions are summarized as follows:
@@
    To address the issue of limited feature diversity in short moments, we propose a novel augmentation strategy based on two types of mixing methods, specifically tailored for video moment retrieval.
@@
    To enhance both center and length prediction for short moments, we introduced length conditioning into DETR-based MR methods for the first time, effectively creating “length-wise expert” queries with length-wise matching.
@@
    Our approach notably enhanced performance on the various MR datasets, resulting in significant improvement of mAP in QVHighlights (9.36%; 41.22 $\rightarrow$ 45.08) and across other datasets.
@@ 2.1 Moment Retrieval
The moment retrieval (MR) task involves identifying the start and end timestamps of a specific moment within a video given a text query. Traditional methods approach MR through either two-stage [2, 13, 45, 47, 9] or one-stage [5, 34, 27, 41, 14, 22, 46] framework, which often rely on cumbersome processes such as proposal generation, non-maximum suppression (NMS), and human priors. A significant paradigm shift was introduced with Moment-DETR [19], which framed MR as a direct set prediction task using Detection Transformer (DETR) [4]. This approach eliminates these dependencies, thereby making MR more efficient and flexible. Subsequent models, including QD-DETR [26], and MESM [25], have further optimized DETR-based MR by addressing cross-modal alignment and refining prediction accuracy. Recent works, such as TR-DETR [32], TaskWeave [44], and UVCOM [38], leverage task-specific features to bridge Moment Retrieval and Highlight Detection, enhancing robustness across both tasks. Additionally, there are researches to develop boundary-aware notations for moments [18] and to explore advanced feature in MR [24, 37]. Considering the efficiency and advancements of DETR-based architectures, our research specifically targets the challenges within DETR-based methods.
Different from the previous works, to the best of our knowledge, we are the first to address the critical limitation of performance degradation on short moments, a common challenge across existing DETR-based methods. Our approach specifically targets the accurate retrieval of short-duration moments, thereby filling a significant gap in the current landscape of DETR-based MR research.
Figure 2: Data perspective analysis. t-SNE [33] visualization of visual features for 50 sampled short moments and 50 non-short moments from QVHighlights train set. Each sampled moment is taken from a distinct video. This plot demonstrates that short moments exhibit fewer and significantly sparser visual features compared to non-short moments, highlighting the limited visual information inherent to their short duration.
@@ 2.2 Mixing-based Augmentation
Mixing-based augmentations have been explored in both image and video tasks, each adapting spatial or temporal mixing based on task requirements. In image classification, Mixup [42] and CutMix [39] create new image samples by interpolating or combining patches, promoting diverse feature representations. Copy-Paste [12] augments data for detection and segmentation by inserting objects from one image into another, increasing object and scene variety. In video understanding tasks, VideoMix [40] inserts randomly selected video cuboid patches from one video into another, thereby introducing both spatial and temporal diversity. Similarly, VIPriors [16] extends traditional image-based mixing augmentations to the temporal dimension, which strengthens temporal feature representations and improves model robustness against temporal fluctuations.
However, these approaches primarily focus on modifying spatial features and are not directly applicable to the Moment Retrieval framework, which relies solely on frame-level features without spatial dimensions. To address this limitation, we introduce a novel data augmentation technique specifically designed for moment retrieval that enhances the diversity of frame-level features.
@@ 3.1 Motivation
Figure 4: Overview. In Section 3.2, we propose MomentMix, which comprises two types of mix-based data augmentations: ForegroundMix and BackgroundMix. ForegroundMix implementation involves cutting a long moment into shorter segments, which are then shuffled to generate new short-moment data. The BackgroundMix preserves the foregrounds unchanged and replaces the backgrounds with temporal segments randomly cropped from other video samples. In Section 3.3, We introduce a Length-Aware Decoder, which utilizes a length-wise bipartite matching process to create length-wise expert decoder queries.
@@
Performance limit of DETR-based methods. Recent approaches actively employ DETR for the MR task and achieve impressive performances. We selected representative DETR-based models and analyzed their performance according to the length of target moments. Despite achieving strong performance, these models exhibited significant performance drops in retrieving short moments, as shown in Table 1. Specifically, QD-DETR, TR-DETR, and UVCOM experienced mAP declines of - 79.8%, 78.0%, and 72.4% for short moments. These results indicate a consistently significant decrease relative to overall performance. To address this degradation, we investigated the underlying causes from both data-centric and model-centric perspectives.
@@
For the data-centric analysis, we examined the statistical characteristics of short moments. As shown in Figure 1, while the total number of short moments is comparable to that of other types of moments, the number of videos containing short moments is clearly limited. This led us to hypothesize that short moments might lack diverse contextual representation and could exhibit a narrow distribution in the training data. To test this hypothesis, we compared the feature distribution of short moments to that of other moments using feature visualization. We randomly sampled 50 short moments and 50 non-short moments from the training set and applied t-SNE [33] to visualize their visual features. As shown in Figure 2, the distribution of visual features for short moments was concentrated, indicating a significant lack of diversity. This observation suggests that the training data for short moments does not capture a wide range of visual features, leading to suboptimal generalization performance during testing.
@@
For the model-centric analysis, we evaluated the model’s prediction tendencies by separately assessing the center and length predictions of short moments compared to other types of moments. As illustrated in Figure 3, only 37% of center predictions for short moments have their center within the ground truth, while 74% for middle moment and 82% for long moments. This revealed that inaccuracies in center prediction are a significant source of overall error.
@@
To overcome these limitations, we propose two novel techniques, MomentMix and Length-Aware Decoder that can be easily integrated into other DETR-based models. Our overall architecture is illustrated in Figure 4, following the design of the common baseline, QD-DETR.
3.2 MomentMix: Leveraging ForegroundMix and BackgroundMix
@@
We propose MomentMix, a data augmentation strategy designed to address the issue of low feature diversity for short moments. MomentMix consists of two components: (1) ForegroundMix, which enhances the diversity of foreground features, and (2) BackgroundMix, which increases the variety of background features. To the best of our knowledge, this is the first data augmentation approach specifically tailored for video moment retrieval.
@@
ForegroundMix augmentation. The goal of ForegroundMix is to increase the visual diversity of foreground features in short moments, enabling more generalized prediction. To achieve this, we randomly extract and mix rich foreground features from longer samples to create augmented short moments. Visual features within a single video naturally exhibit higher similarity compared to those from different videos. By exploiting features from other video clips, our method allows the model to generalize diverse contexts, ensuring it can detect short moments reliably even in challenging or ambiguous frames.
@@
These sub-regions represent segments of the foreground, uniformly sampled as $f_{i}^{s},f_{i}^{e}\sim\text{Unif}(s,e)$, where $s$ and $e$ indicate start and end of $f_{\text{source}}$. Similarly, the background region, $b_{\text{source}}=b_{\text{front}}\cup b_{\text{back}}$, is divided into $n+1$ sub-regions denoted as $b_{0},b_{1},\dots,b_{n}$, representing the segments of the background as:
@@
$\begin{gathered}\pi:\{f_{1},f_{2},\dots,f_{n}\}\rightarrow\{f_{1}^{\prime},f_{2}^{\prime},\dots,f_{n}^{\prime}\},\\ \pi:\{b_{0},b_{1},\dots,b_{n}\}\rightarrow\{b_{0}^{\prime},b_{1}^{\prime},\dots,b_{n}^{\prime}\},\\ \end{gathered}$ (3)
where $\pi$ be a random permutation function.
Each shuffled foregrounds $\{f^{\prime}_{i}\}_{i=1}^{n}$ is then paired with backgrounds $\{b^{\prime}_{i}\}_{i=0}^{n}$ to form the following augmented samples:
@@
BackgroundMix augmentation. The goal of BackgroundMix is to improve the diversity of the visual background features, thereby strengthening the association between foreground visual features and the text query. To achieve this, we keep the original foreground features while replacing the background with features from different videos. This method provides the model with richer training signals, allowing it to learn various boundaries more effectively.

### v2

# MomentMix Augmentation with Length-Aware DETR for Temporally Robust Moment Retrieval

@@ Abstract
Video Moment Retrieval (MR) aims to localize moments within a video based on a given natural language query. Given the prevalent use of platforms like YouTube for information retrieval, the demand for MR techniques is significantly growing. Recent DETR-based models have made notable advances in performance but still struggle with accurately localizing short moments. Through data analysis, we identified limited feature diversity in short moments, which motivated the development of MomentMix. MomentMix generates new short-moment samples by employing two augmentation strategies: ForegroundMix and BackgroundMix, each enhancing the ability to understand the query-relevant and irrelevant frames, respectively. Additionally, our analysis of prediction bias revealed that short moments particularly struggle with accurately predicting their center positions and length of moments. To address this, we propose a Length-Aware Decoder, which conditions length through a novel bipartite matching process. Our extensive studies demonstrate the efficacy of our length-aware approach, especially in localizing short moments, leading to improved overall performance. Our method surpasses state-of-the-art DETR-based methods on benchmark datasets, achieving the highest R1 and mAP on QVHighlights and the highest R1@0.7 on TACoS and Charades-STA (such as a 9.62% gain in R1@0.7 and an 16.9% gain in mAP average for QVHighlights). The code is available at https://github.com/sjpark5800/LA-DETR.
@@ 1 Introduction
As vast amounts of video content are created and shared on the internet daily [5], moment retrieval (MR) [1, 9] has gained significant attention to improve user experience and search efficiency. MR identifies the specific moments within a video that best align with a given query. Specifically, this task involves localizing the start and end points in the video relevant to the textual query, offering a more detailed understanding of video content.
@@
Figure 1: (a) Performance (mAP) of previous moment retrieval (MR) DETR-based methods [25, 31, 36] on QVHighlights test set by the lengths of the moment. Notice that the mAP drops significantly in capturing short-moment, where MR can be best utilized. (b) Short moment performance (mAP) comparison.
For the MR task, various methods have been explored, including DETR-based approaches [17, 25, 24, 31, 36], non-DETR models [19, 22, 2], and models leveraging the power of LLMs [29, 11, 13]. LLM-based approaches achieve strong zero-shot performance but are impractical for real-world applications due to slow inference speeds. Non-DETR methods offer faster inference but rely heavily on complex post-processing, such as Non-Maximum Suppression (NMS), making them sensitive to IoU thresholds. DETR-based models have attracted significant attention by addressing these issues through their fast inference speeds and robust set-prediction framework, enabling accurate predictions without additional post-processing. However, our empirical findings indicate that these DETR-based models suffer from a significant drop in performance when handling short moments as highlighted in Figure 1. For example, UVCOM [36] shows an average mAP of 44.90 for middle-length moments (10–30 seconds), while achieving only 10.67 for short moments (less than 10 seconds), revealing a substantial gap.
@@
Retrieving short moments within videos is a crucial task because videos often contain a significant amount of redundant or irrelevant information, while essential content is frequently condensed in short moments. This aligns with the importance of MR, where improving the accuracy of short-moment retrieval enables the precise extraction of the most relevant information. Such improvements can significantly reduce the time and effort required for video exploration. For example, highlights in sports and news, as well as key scenes in movies and dramas, often involve short moments. This emphasizes the importance of accurately retrieving short moments in practical scenarios.
@@
In this study, we analyzed the challenges associated with existing methods in short-moment retrieval from both data and model perspectives. From a data perspective, we examined the feature distribution of short moments compared to other moments. As shown in Figure 2, the features of short moments tend to be more concentrated around the mean feature, with 42.9% of samples falling within one standard deviation. In contrast, for other moments, 26.6% of samples fall within one standard deviation of the mean. These observations show that short moments exhibit relatively simple and less diverse feature distributions. On the model side, we analyzed the trends in prediction accuracy by breaking down the model’s final output into center and length components, as illustrated in Figure 3. Although conceptually, moments are defined by (start, end), existing models predict them in the format of (center, length). Interestingly, we found that the accuracy of both center and length predictions for short moments was significantly lower than for other moment types.
@@
Based on the above analysis, we propose a new DETR-based MR framework that addresses the performance degradation when retrieving short moments. Our framework consists of two novel techniques: a data augmentation technique called MomentMix and a Length-Aware Decoder (LAD). Through data analysis, we identified a key limitation in the feature diversity of short moments. It prompts us to design MomentMix, a two-stage mix-based augmentation strategy specifically to enhance the diversity of short-moment samples. In a video sample, we define the temporal moments relevant to the text query as the foreground and the unrelated moments as the background. MomentMix operates in two stages. In the first stage, we create new short foregrounds from rich foreground elements by cutting and mixing. In the next stage, to further diversify these short samples, we utilize portions of other videos as backgrounds, forming varied foreground-background combinations. This increased feature diversity of short moments leads to more robust detection of short moments (see Fig. A1 in Supple.).
@@
Our analysis of model outputs revealed that both center and length prediction errors significantly contribute to performance drops in short-moment retrieval. To address this, we introduce a length-aware decoder designed to enhance center prediction by conditioning it on length. Specifically, we predefine length classes (e.g., short, middle, long) and uniformly assign each decoder query to these length-specific classes. Additionally, we modify the bipartite matching process so that queries are matched with ground-truth moments within the same length class. This approach improves the accuracy of both center and length predictions for short moments.
Our contributions are summarized as follows:
@@
    To address the issue of limited feature diversity in short moments, we propose a novel two-stage mix-based augmentation strategy, specifically tailored for short moment retrieval.
@@
    To enhance center predictions for short moments, we introduced length conditioning into DETR-based MR methods for the first time, effectively creating “length-wise expert” queries with length-wise matching.
@@
    Our approach notably enhanced performance on the various MR datasets, resulting in significant improvement of mAP in QVHighlights (16.9%; 39.84 $\rightarrow$ 46.61) and across other datasets.
@@ 2.1 Moment Retrieval
Moment retrieval (MR) has been explored through DETR-based, non-DETR, and LLM-based approaches. DETR-based methods [17] introduced a paradigm shift by formulating MR as a set prediction task, eliminating proposal dependency and extensive post-processing. Recent works [25, 24, 23, 42, 31, 36] further enhance video-text alignment.
Despite the advantages of DETR-based methods, recent non-DETR approaches [19, 21, 2] have introduced unified frameworks that integrate additional video temporal understanding tasks, including video summarization and highlight detection. While these methods achieve strong performance, they rely heavily on handcrafted heuristics such as Non-Maximum Suppression (NMS), making them computationally expensive and sensitive to hyperparameters. LLM-based methods [29, 11, 13] leverage large-scale pre-trained models for improved textual understanding and reasoning. However, their slow inference and high computational cost hinder their practicality in real-time applications. Given the trade-offs across approaches, DETR-based models remain the practical solution for MR, balancing accuracy, efficiency, and scalability.
Recent findings from FlashVTG [2] indicate performance degradation in short moments within the MR task. To the best of our knowledge, our study is the first to identify and analyze the underlying causes of this issue from both model and data perspectives, proposing novel approaches tailored for short moment performance improvement.
Figure 2: Data Perspective Analysis. t-SNE [32] visualization of visual features distribution for 50 sampled short moments and 50 non-short moments from QVHighlights train set. Each sampled moment is taken from a distinct video. This plot demonstrates that short moments exhibit fewer and significantly sparser visual features compared to non-short moments, highlighting the limited visual information inherent to their short duration.
@@ 2.2 Mixing-based Augmentation
Mixing-based augmentations have been explored in both image and video tasks, each adapting spatial or temporal mixing based on task requirements. In image classification, Mixup [40] and CutMix [37] create new image samples by interpolating or combining patches, promoting diverse feature representations. Copy-Paste [10] augments data for detection and segmentation by inserting objects from one image into another, increasing object and scene variety. In video understanding tasks, VideoMix [38] inserts randomly selected video cuboid patches from one video into another, thereby introducing both spatial and temporal diversity. Similarly, VIPriors [14] extends traditional image-based mixing augmentations to the temporal dimension, which strengthens temporal feature representations and improves model robustness against temporal fluctuations. However, these approaches primarily focus on modifying spatial features and are not directly applicable to the Moment Retrieval framework, which relies solely on frame-level features without spatial dimensions. A related augmentation [20] has been proposed for long-video moment retrieval, employing proposal-based methods by varying proposal lengths and inserting proposals into other videos. In contrast, our work explicitly targets the lack of feature diversity in short moments within a proposal-free DETR-based framework. To the best of our knowledge, we are the first to propose an augmentation strategy specifically tailored to short-moment retrieval in proposal-free methods.
@@ 3.1 Motivation
Why DETR-based Methods Underperform on Short Moments? Recent approaches actively employ DETR [3] for the MR task because of its impressive performance and its end-to-end design that eliminates the need for extra post-processing. We selected representative DETR-based models and analyzed their performance according to the length of target moments. Despite achieving strong performance, these models exhibited significant performance drops in retrieving short moments. Specifically, for short moments, QD-DETR, TR-DETR, and UVCOM experienced mAP declines of 79.8%, 78.0%, and 72.4%, respectively, compared to their overall performance. To address this degradation, we investigated the underlying causes from both data-centric and model-centric perspectives.
@@
For the data-centric analysis, we examined the statistical characteristics of short moments. As illustrated in QVHighlights [17], while the total number of short moments is comparable to that of other types of moments, the number of videos containing short moments is clearly limited. This led us to hypothesize that short moments might lack diverse contextual representation and could exhibit a narrow distribution in the training data. To test this hypothesis, we compared the feature distribution of short moments to that of other moments using feature visualization. We randomly sampled 50 short moments and 50 non-short moments from the training set and applied t-SNE [32] to visualize the distribution of their visual features. As shown in Figure 2, the distribution of visual features for short moments was concentrated, indicating a significant lack of diversity. This observation suggests that the training data for short moments does not capture a wide range of visual features, leading to suboptimal generalization performance during testing.
@@
For the model-centric analysis, we evaluated the model’s prediction tendencies by separately assessing the center and length predictions of short moments compared to other types of moments. As illustrated in Figure 3, only 37% of center predictions for short moments have their center within the ground truth, while 74% for middle moments and 82% for long moments. This revealed that inaccuracies in center prediction are a significant source of overall error.
@@
To overcome these limitations, we propose two novel techniques, MomentMix and Length-Aware Decoder that can be easily integrated into other DETR-based models. Our overall architecture follows the design of QD-DETR [25].
3.2 MomentMix: Augmentation for Short Moment
Figure 4: MomentMix. A two-stage mix-based data augmentation technique for generating short-moment samples. The first stage, ForegroundMix, splits a long moment into shorter clips and shuffles them. The second stage, BackgroundMix, preserves the foreground while replacing the background with randomly cropped temporal segments from other videos.
@@ 3.2 MomentMix: Augmentation for Short Moment
We propose MomentMix, a novel two-stage data augmentation strategy specifically designed to address the issue of low feature diversity for short moments. In the first stage (ForegroundMix), diverse short foregrounds are generated by cutting and mixing longer foreground samples. In the second stage (BackgroundMix), background diversity for augmented short samples is further enhanced by recombining backgrounds from various video clips. To the best of our knowledge, this is the first data augmentation approach specifically tailored for robust short moment retrieval.
@@
First Stage: ForegroundMix. The goal of ForegroundMix is to increase the visual diversity of foreground features in short moments, enabling more generalized prediction. To achieve this, we randomly extract and mix rich foreground features from longer samples to create augmented short moments. Visual features within a single video naturally exhibit higher similarity compared to those from different videos. By exploiting features from other video clips, our method allows the model to generalize diverse contexts, enabling it to leverage a broader range of visual features for robust short-moment detection.
@@
These sub-regions represent segments of the foreground, uniformly sampled as $f_{i}^{s},f_{i}^{e}\sim\text{Unif}(s,e)$, where $s$ and $e$ indicate the start and end of $f_{\text{source}}$. Similarly, the background region, $b_{\text{source}}=b_{\text{front}}\cup b_{\text{back}}$, is divided into $n+1$ sub-regions denoted as $b_{0},b_{1},\dots,b_{n}$, representing the segments of the background as:
@@
[…]
where $\pi$ be a random permutation function. Each shuffled foregrounds $\{f^{\prime}_{i}\}_{i=1}^{n}$ is then paired with backgrounds $\{b^{\prime}_{i}\}_{i=0}^{n}$ to form the following augmented samples:
@@
To prevent temporal disruption during temporal mixing, we exclude queries involving explicit temporal relations from augmentation through preprocessing.
Second Stage: BackgroundMix. The goal of BackgroundMix is to improve the diversity of the visual background features of augmented short samples from ForegroundMix, thereby strengthening the association between foreground visual features and the text query. To achieve this, we keep the original foreground features while replacing the background with features from various videos. This method provides the model with richer training signals, allowing it to learn various boundaries more effectively for robust short moment retrieval.